03-Consider the product structure tree shown below: D(2) C(3) A(2) E(3) B(2) E(3) A(3) With a beginning inventory of zero for all components and finished items, how many assembly Es are needed to produce 25 product Xs? A) 125 B) 375 C) 400 D) 250 Q4-Livernonions Industries uses a Kanban system to manage production of the BR-549 subassembly. In an eight-hour day, they use 900 units and use containers that can hold 30 units each. It takes 120 minutes to fill and move a container of parts to the right place on the BR-549 line that is currently rated at a 15% safety factor. How many production hours of inventory are needed? A) 7.6 B) 4.3 C) 18.4 D) 2.4
Answer #1

1. Option B 375 E's

1. Produce one product of X we need 2 product of D and 3 products of C. Each product of D required 3 materials of E and Each Product of C required 3 Materials of E

One Product of X required = 3 * 3 + 2 * 3 = 15 Metrails of E

Thus to Produce 25 Product of X = 25 * 15 = 375 E's

2. Option D. 2.4 Hours

Number of Hours in a day * Safety Factor * Number of Hours to Fill and Use

8 * 15% * 2

= 2.4 Hours